OpenStudy (anonymous):

solve using substituion method 2x+5y=-5 x+3y=3

OpenStudy (anonymous):

I don't think you can divide by two and then subtract randomly off of the other parts. What is the direct question?

OpenStudy (anonymous):

Well It just says "solve using substitution or elimination methods. I have 3 questions like this.

OpenStudy (anonymous):

No, the actual question please. Or are those two questions? It looks like you divided 2x by 2 to get just x but only subtracted 2 from the other ones. I don't think -5 divided by 2 is 3. Or is it? Correct me if I'm wrong from some crazy formula but yeah.

OpenStudy (anonymous):

its asking me to solve that system using whatever method. I dont understand which method to use or how to do it. im not sure what your thinking lol.

OpenStudy (anonymous):

2x+5y=-5 x+3y=3 2 (x + 3y = 3) = 2x + 6y = 6 2x + 5y = -5 2x + 6y = 6 Subtract the bottom equation from the top. 2x + 5y = -5 -2x - 6y = -6 -y = -11 Divide both sides by -1. y = 11 2x + 5(11) = -5 2x + 55 = -5 2x = -60 x = -30 2x + 6(11) = 6 2x + 66 = 6 2x = -60 x = -30 y = 11 and x = -30

OpenStudy (anonymous):

I appreciate it. I am not good at math at all